OpenStudy (anonymous):

Please Help! A solution is made by dissolving 15.5 grams of glucose (C6H12O6) in 245 grams of water. What is the freezing-point depression of the solvent if the freezing point constant is -1.86 °C/m? Show all of the work needed to solve this problem.

OpenStudy (aaronq):

use the formula: \(\Delta T=i*m*K_{f}\) More on this: http://openstudy.com/study#/updates/521ecb3ae4b0750826e0c362

OpenStudy (anonymous):

Thanks!

OpenStudy (aaronq):

no problem !

OpenStudy (anonymous):

OK So I looked at the other problem but im still confused. could you walk me through the steps with this problem?

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

sure. first you have to find the molality, can you do that? \(molality=\dfrac{moles ~of~solute}{kg~of ~solvent}\)

OpenStudy (anonymous):

What part is the Molaity? i know it shold be \[\frac{ 15.5 }{ 245 }\]

OpenStudy (anonymous):

does that make the molaity =.0632?

OpenStudy (aaronq):

molality is m. you have to convert the mass of glucose to moles first, as well as converting g to kg.

OpenStudy (anonymous):

oh ok \[\frac{ .086 m}{ .245kg }\] Is this right

OpenStudy (anonymous):

and that = .35?

OpenStudy (aaronq):

yeah, except 0.086 is in moles, not m now you can just multiply it by the \(K_f\), to get \(\Delta T \).

OpenStudy (anonymous):

\[DeltaT= (1)(0.35)(1.86)=.65\]

OpenStudy (anonymous):

so the freezing point would be -.65?

OpenStudy (anonymous):

can you help me with another question?

OpenStudy (anonymous):

A solution is made by dissolving 2.5 moles of sodium chloride (NaCl) in 198 grams of water. If the molal boiling point constant for water (Kb) is 0.51 °C/m, what would be the boiling point of this solution? Show all of the work needed to solve this problem.

OpenStudy (aaronq):

yep, because the normal freezing point is zero celsius, it is now -0.65 celsius

OpenStudy (aaronq):

This question is very similar to the link i posted, nonetheless i can help you through it. First find the molality

OpenStudy (anonymous):

\[\frac{ 2.5M }{ .198kg }\]

OpenStudy (anonymous):

12.62m

OpenStudy (aaronq):

now multiply it by \(K_b\). Then you have to determine \(i\), the van't hoff constant.

OpenStudy (aaronq):

This is the number of particles (whether it's ions or polyatomic ions) that the substance separates into when dissolved in water. so, you kinda have to know a little about their behaviour when in water.

OpenStudy (anonymous):

and the kb is .51?

OpenStudy (aaronq):

yes

OpenStudy (aaronq):

So for i, if you had \(MgCl_2\rightarrow, Mg^{2+}+2Cl^-\), i=3

OpenStudy (anonymous):

So its 6.4 but im a little confused now. what is the equation above?

OpenStudy (aaronq):

it's an example of how to determine what number for \(i\) you need to use

OpenStudy (anonymous):

Oh ok so is the answer 106.4C?

OpenStudy (aaronq):

what value for \(i\) did you use?